# Aligning Machine and Human Visual Representations across Abstraction Levels

### v3

@@ A.1.2 Data generation
Image data. For creating AligNet, we use the publicly available ImageNet database [Deng et al., 2009]. ImagNet is a natural image dataset with $\approx 10^{6}$ training data points and $1000$ image categories [Russakovsky et al., 2015]. The categories are almost equally distributed in the data with small variations in the number of images between the different classes. Hence, ImageNet can be considered a highly balanced dataset. ImageNet has been the dominant image dataset for training large computer vision models until the advent of image/text multi-modal training a few years ago. Although to date larger image datasets exists, ImageNet is still is one of the largest open-source and most widely used image datasets in the field of Computer Vision.
@@
Random. Uniform random sampling is the vanilla sampling approach and was used to create the THINGS datasets (see Sec. A.1.1). In random sampling, three images are chosen uniformly at random without replacement from all of the $m$ images in the data to create a triplet. Since there are $C=1000$ classes and each class has approximately $1000$ images, most of the triplets generated with this approach contain three images from three different classes. The number of triplets different from triplets with images from three distinct classes is negligible. Note that this is the same sampling approach that was used to generate the THINGS triplets [Hebart et al., 2020]. A triplet generated via random sampling can be defined as the following triplet set $\mathcal{S}\coloneqq\{x_{i},x_{j},x_{k}\}$ with the constraint $\left(y_{i}\neq y_{j}\neq y_{k}\right)$.
@@ A.1.3 Objective function
$\mathcal{L}_{\mathrm{alignet}}({\bm{S}}^{\prime},{\bm{S}}^{\dagger})\coloneqq\frac{1}{B}\sum_{s=1}^{B}\log\sigma\left(\left[S^{\prime}_{i,j},S^{\prime}_{i,j},S^{\prime}_{j,k}\right],\tau^{\prime}\right)_{s}-\sigma\left(\left[S^{\prime}_{i,j},S^{\prime}_{i,j},S^{\prime}_{j,k}\right],\tau^{\prime}\right)_{s}\log\sigma\left(\left[S^{\dagger}_{i,j},S^{\dagger}_{i,k},S^{\dagger}_{j,k}\right],\tau^{\dagger}\right)_{s},$ (5)
@@ A.1.4 Surrogate teacher model
Muttenthaler et al. [2023a] showed that image/text models and models trained on large, diverse datasets are better aligned with human similarity judgments than vision models trained with a self-supervised learning objective or supervised models trained on ImageNet. Thus, we use the best-performing image/text model according to various computer vision benchmarks at the time of writing this manuscript as our teacher model. This model is called SigLIP [Zhai et al., 2023]. SigLIP, similar to CLIP [Radford et al., 2021] and ALIGN [Jia et al., 2021], is trained via contrastive language-image pretraining using millions of image/text pairs. The difference between CLIP and SigLIP is that SigLIP uses a paired sigmoid loss instead of the standard $\mathrm{softmax}$ function usually used for pretraining image/text models via cross-entropy. Image/text pretraining allows the model to learn an aligned representation space for images and text; thus, adding more semantic information about the objects in an image to the model representations.
We use the SigLIP-So400m variant of SigLIP as our teacher model. The So400m variant uses an optimized ViT backbone whose performance is similar to one of the largest ViTs, ViT-g/14 [Alabdulmohsin et al., 2023] while having fewer parameters and thus being smaller. The number of parameters of SoViT-400m/14 lies somewhere between ViT-L/16 and ViT-g/14. The output dimensionality of the image and text encoder representations of SoViT-400m/14 is $p$=$1152$ each. We align the image encoder representations with human odd-one-out choices using the UD optimization outlined in Eq. 6. This allows us to increase the triplet odd-one-out accuracy of SigLIP-So400m from $44.24\%$ to $61.7\%$ (see rightmost column in Tab. 1) which is close to the human-noise ceiling of $66.67\%$ for THINGS [cf. Hebart et al., 2020] and thus among the best human-aligned models without AligNet fine-tuning [cf. Muttenthaler et al., 2023a]. Note that this is a relative increase in performance of $39.47\%$. Throughout this manuscript we use the human-aligned version of SigLIP-So400m as the surrogate teacher model for generating human-like similarity judgments and distilling human-like similarity structure into student VFMs. We select a diverse and representative set of student VFMs.
@@ A.1.5 Student Models
Since previous research has demonstrated that a model’s architecture has no significant impact on the degree of alignment with human similarity judgments [Muttenthaler et al., 2023a, b], we use the same architecture for all student models that we fine-tune on AligNet. Specifically, we use the Vision Transformer [ViT; Dosovitskiy et al., 2021] for the backbone of each student model. We use the ViT rather than CNN-based model because ViTs have recently emerged as the dominant neural network architecture for computer vision application and VFMs. Every large VFM that is used in practice is based on the ViT [Tschannen et al., 2023; Chen et al., 2023; Zhai et al., 2023; Beyer et al., 2024]. Unless otherwise mentioned, we use the base model size, i.e., ViT-B. ViT-B has $12$ attention layers and an internal (hidden) representation size of $p=768$. It has been shown that both the training data and the objective function have a substantial impact on the degree of alignment with human behavior. Thus, we use student models that were trained on different pretraining task with different training data and objective functions.
Supervised pre-training is still the prevailing mode of training computer vision models. Therefore, we trained ViT-B on the popular ImageNet dataset consisting of 1.4M natural images [Deng et al., 2009]. To examine how model performance changes as a function of the size of a model, we train ViT instances of three different sizes on ImageNet: ViT-S/16, ViT-B/16, ViT-L/16. The image patch size is the same for each of those models. To evaluate the effect of AligNet on self-supervised pretraining, we use pretrained DINOv1 [Caron et al., 2021] and DINOv2 [Oquab et al., 2024] models of which DINOv1 was pretrained on ImageNet and DINOv2 was pretrained on a different, larger image dataset as denoted below. In addition, we investigate multimodal training of vision models that add textual information both in the form of image captioning via the CapPa model [Tschannen et al., 2023], as well as contrastive language-image pretraining (CLIP) via SigLIP [Zhai et al., 2023]. The latter model is considered state-of-the-art on many downstream computer vision applications and is used as the image embedding model in modern large visual-language models [VLMs; Chen et al., 2023; Beyer et al., 2024]. The full list of student models that we consider in our analyses is the following:
@@
    • Training data: ImageNet [Deng et al., 2009]
@@
  • SigLIP-{B,SO400M}
    • Training data: WebLI [Zhai et al., 2023]
    • Objective: CLIP [Radford et al., 2021]
@@
    • Training data: ImageNet [Deng et al., 2009]
    • Objective: Self-supervised image pretraining [Caron et al., 2021]
@@
    • Objective: Self-supervised teacher-student distillation [Oquab et al., 2024]
@@
    • Objective: Multi-modal image captioning [Tschannen et al., 2023]
@@ A.2.1 Multi-arrangement task
Human similarity judgments for King et al. [2019] and Cichy et al. [2019] were obtained by using a multi-arrangement task. In a multi-arrangement task, participants are presented with a computer screen showing images of several different objects. The participants are asked to arrange the images into semantically meaningful clusters, given the instruction that images of objects that lie close together are considered more similar. From this arrangement, one can infer pairwise (dis-)similarities of the objects and average those across all participants to obtain a representative (dis-)similarity matrix.
@@ A.2.3 Neural network representations
Representation Similarity Matrices (RSMs) for neural network representations are obtained by first embedding the same set of images that were presented to the human participants in the $p$-dimensional latent space of a model. The latent space could be any layer of a neural network. For the base models, we use the representations of the image encoder for SigLIP and the CLS token of the penultimate layer for CapPa, DINOv2, and ViT-B. We do this because previous work has shown that the penultimate layer space and the image encoder space of image/text models respectively yield the highest similarity to human behavior [Peterson et al., 2018, 2019; Muttenthaler et al., 2023a].
@@
$\leavevmode\nobreak\ \text{cos}({\bm{x}}_{i},{\bm{x}}_{j})\coloneqq\frac{{\bm{x}}_{i}^{\top}{\bm{x}}_{j}}{||{\bm{x}}_{i}||_{2}||{\bm{x}}_{j}||_{2}};$ $\leavevmode\nobreak\ \phi({\bm{x}}_{i},{\bm{x}}_{j})\coloneqq\frac{\left({\bm{x}}_{i}-\bar{{\bm{x}}_{i}}\right)^{\top}\left({\bm{x}}_{j}-\bar{{\bm{x}}_{j}}\right)}{||\left({\bm{x}}_{i}-\bar{{\bm{x}}_{i}}\right)||_{2}||\left({\bm{x}}_{j}-\bar{{\bm{x}}_{j}}\right)||_{2}},$
@@ A.3 Levels data
Stimuli. The experimental stimuli were images taken from the ImageNet dataset [Deng et al., 2009]. Another 9 images were used for instructions only and depicted natural objects selected from the Bank of Standardized Stimuli [BOSS; Brodeur et al., 2014], available at BOSS. We grouped the visual stimuli presented in the triplets according to different levels of abstraction: coarse-grained semantic, which comprised three images from three different categories; fine-grained semantic, showing three images from the same category; and class-boundary, where two images were from the same and one from a different category.
Instead of randomly sampling triplets—which would reproduce dataset biases—we stratified sampling by superclasses. ImageNet classes follow the WordNet hierarchy Deng et al. [2009]; Russakovsky et al. [2015], which includes higher-level classes. For instance, all dog breeds can be summarized as a single dog superclass. To avoid presenting dogs, birds, and other fine-grained classes that are overrepresented in ImageNet more frequently to the participants than other categories, we grouped the ImageNet classes into $717$ coarse-grained WordNet superclasses. We uniformly at random sampled images from those $717$ superclasses to construct the different kinds of triplets. Note that for all superclasses with more than one class, we uniformly at random chose one subclass and either uniformly at random sampled one image, two images (without replacement), or three images (without replacement) from that subclass, depending on the triplet type. For most superclasses that were comprised by a single subclass only, i.e., a one-to-one-mapping, we could skip the subclass sampling part. Triplet sampling resulted in ($N=450$) predefined experiment trial sets, of which ($N=448$) were used for testing. Across these, each triplet was presented within ($N=5$) different experiment files. This sampling process ensured a balanced distribution of triplets across the sample, and the repetition of each triplet in five different participants allowed for the calculation of an uncertainty distribution for each triplet.
@@
Procedure. The experiment was run online using jsPsych v7.3.3 and custom plugins. Participants were asked to provide demographic information, including their age and gender. Thereafter, they viewed written instructions about the task and performed six practice trials (2 trials per triplet level of abstraction). Participants were free to repeat the instructions until they felt confident to perform the experiment. The experiment proper comprised $N=330$ experiment trials. Each trial started with a fixation cross (1s), followed by the presentation of a triplet (max. 15s). Participants were asked to select the “odd-one-out” using the right, left, or downward facing arrow keys on their keyboard. Responses could be entered between 1-15s after triplet onset, after which the next trial started. Trials in which participants failed to submit a response were rare (M = 0.27% of trials; min = 0.00%, max = 6.06%). The serial order of triplet types (e.g., fine-grained or coarse-grained semantic) and ImageNet classes (e.g., dogs or birds) was counterbalanced across the experiment. We additionally counterbalanced the serial position of trial types across participants using a Latin-Square approach [LSD; Grant, 1948]. Participants could take short breaks (self-paced for up to 2 minutes) after $N=50,150,$ and, $200$ experiment trials; Experimental trials were interleaved with $N=16$ catch trials (class border triplets), which were predefined based on low model uncertainty and 100% agreement among participants on these specific triplets during piloting. Catch trial performance was used as an indicator of adequate task engagement (see participant inclusion criteria above).
@@
Appendix B Additional/Detailed Results
In addition to the results that we presented in the main paper, here, we report the same set of results in more detail and add further analyses that we touched upon in the main part but did not go into much detail. We start by presenting findings from a comprehensive set of evaluations that we performed to test the alignment of model representations with human similarity judgments both for similarity judgments collected in previous efforts [cf. Peterson et al., 2018; Cichy et al., 2019; King et al., 2019; Hebart et al., 2020] and for our own similarity judgment dataset Levels (see Section A.3). Subsequently, we demonstrate benefits of our method for machine learning downstream tasks that test generalization, such as few-shot learning and out-of-distribution detection. Finally, we perform a rigorous qualitative analysis of the changes in the model representations after applying our alignment framework to the various student models.
@@ B.1 Human Alignment
  • Distillation without alignment: We perform AligNet fine-tuning without the first step (see Fig. 1) of aligning the teacher model with the human similarity judgments in THINGS. That is, we distill the teacher similarity structure of the pretrained, non-aligned SigLIP-So400m model (see Sec. A.1.4) into a student VFM.
@@ B.1.1 Detailed results for RSA
In Tab. 1 we present the average alignment scores for all student models that we considered in our analyses (see Sec. A.1.5) and each of the four representation settings defined above for the datasets from Cichy et al. [2019], King et al. [2019], and Hebart et al. [2020]. For the former two datasets, alignment is measured via RSA using Spearman rank correlation between the upper triangulars of the pairwise similarity matrices. For those datasets, human similarity judgments were collected using a multi-arrangment task where participants were asked to arrange (natural) objects on a computer screen (see Sec. A.2.1 for details). For the THINGS dataset, alignment is measured via triplet odd-one-out accuracy, i.e., the fraction of triplets for which models selected the same odd-one-out object as the human participants.
@@
ViT-L $0.041$ $\mathbf{0.540}$ $0.227$ $\mathbf{0.751}^{\dagger}$ $0.095$ $0.463$ $0.201$ $0.664$ $34.84\%$ $49.30\%$ $44.40\%$ $60.75\%$
@@
DINOv2 ViT-B $\mathbf{0.492}$ $0.522$ $0.710$ $0.743$ $\mathbf{0.464}$ $0.445$_\__ $0.661$ $\mathbf{51.24}\%$ $49.49\%$ $60.51$% $62.13\%$
@@
SigLIP So400m (Teacher) $0.226$ $0.723$ $0.267$ $\mathbf{0.638}$ $44.24\%$ $\mathbf{61.70}\%$
@@
We find that base models—that is the original representation setting (see above)—performed poorly for almost every human similarity judgment dataset (see Tab. 1 and Fig. 6). The differences among the models in this setting is significant. While DINOv2 ViT-B achieves a Spearman rank correlation coefficient of $\rho=0.492,p<0.001$ for the dataset in Cichy et al. [2019] and an odd-one-out accuracy of $51.21\%$ for the THINGS dataset, ViT-L is not correlated with the human similarity judgments from Cichy et al. [2019] and King et al. [2019] ($\rho=0.041$ and $\rho=0.095$ respectively) and shows a close to chance-level odd-one-out accuracy of $34.84\%$ for the THINGS dataset (see Tab. 1 and $x$-axis in Fig. 7). AligNet fine-tuning significantly improved the alignment scores of all models for all datasets. In addition, it minimized the differences between the models to an extent that their differences in the degree of alignment are not statistically significant. Thus, AligNet fine-tuning models are equally well aligned with human similarity judgments irrespective of their architecture, pretraining data, and objective function.
In the top row of Fig. 6 we show the performance of the transformed representations as a function of the base model performance for those datasets. We can see that AligNet fine-tuning significantly improves upon the base model representations ($x$-axis). It yields the most human-aligned models across all three datasets (depicted by the circles) and often substantially improves upon the linearly aligned representations (depicted by the squares).
@@
In the bottom two rows of Fig. 6 and Fig. 7 respectively we show model performances individually for the single category pairwise similarity rating datasets. We observe that AligNet fine-tuning significantly improved the degree of alignment for all models for the various, animals, and automobiles categories but there does not appear to be a difference in performance between the different model transformation settings for the categories fruits, furniture, and vegetables. We hypothesize that this difference in alignment benefits stems from the fact that animal- and automobile-related concepts are more frequently represented in the THINGS dataset than fruits-, furniture-, and vegetable-related concepts [cf. Hebart et al., 2019, 2020]. Moreover, THINGS is a human similarity judgment dataset that reflects global coarse-grained semantic, similar to the coarse-grained setting in Levels. Thus, it is rather surprising to find benefits for human local fine-grained semantic by either UD or fine-tuning on AligNet.

### v4

@@ A.1.2 Data generation
Image data. For creating AligNet, we use the publicly available ImageNet database [16]. ImagNet is a natural image dataset with $\approx 10^{6}$ training data points and $1000$ image categories [73]. The categories are almost equally distributed in the data with small variations in the number of images between the different classes. Hence, ImageNet can be considered a highly balanced dataset. ImageNet has been the dominant image dataset for training large computer vision models until the advent of image/text multi-modal training a few years ago. Although to date larger image datasets exists, ImageNet is still is one of the largest open-source and most widely used image datasets in the field of Computer Vision.
@@
Random. Uniform random sampling is the vanilla sampling approach and was used to create the THINGS datasets (see Sec. A.1.1). In random sampling, three images are chosen uniformly at random without replacement from all of the $m$ images in the data to create a triplet. Since there are $C=1000$ classes and each class has approximately $1000$ images, most of the triplets generated with this approach contain three images from three different classes. The number of triplets different from triplets with images from three distinct classes is negligible. Note that this is the same sampling approach that was used to generate the THINGS triplets [28]. A triplet generated via random sampling can be defined as the following triplet set $\mathcal{S}\coloneqq\{x_{i},x_{j},x_{k}\}$ with the constraint $\left(y_{i}\neq y_{j}\neq y_{k}\right)$.
@@ A.1.3 Objective function
$\mathcal{L}_{\text{alignet}}({\bm{S}}^{\prime},{\bm{S}}^{\dagger})\coloneqq\frac{1}{B}\sum_{s=1}^{B}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\sigma\left(\left[S^{\prime}_{i,j},S^{\prime}_{i,k},S^{\prime}_{j,k}\right],\tau^{\prime}\right)_{s}}}\log\sigma\left(\left[S^{\prime}_{i,j},S^{\prime}_{i,k},S^{\prime}_{j,k}\right],\tau^{\prime}\right)_{s}$-
$\sigma\left(\left[S^{\prime}_{i,j},S^{\prime}_{i,k},S^{\prime}_{j,k}\right],\tau^{\prime}\right)_{s}\log\sigma\left(\left[S^{\dagger}_{i,j},S^{\dagger}_{i,k},S^{\dagger}_{j,k}\right],\tau^{\dagger}\right)_{s},$ (5)
@@ A.1.4 Surrogate teacher model
Muttenthaler et al. [55] showed that image/text models and models trained on large, diverse datasets are better aligned with human similarity judgments than vision models trained with a self-supervised learning objective or supervised models trained on ImageNet. Thus, we use the best-performing image/text model according to various computer vision benchmarks at the time of writing this manuscript as our teacher model. This model is called SigLIP [86]. SigLIP, similar to CLIP [66] and ALIGN [34], is trained via contrastive language-image pretraining using millions of image/text pairs. The difference between CLIP and SigLIP is that SigLIP uses a paired sigmoid loss instead of the standard $\mathrm{softmax}$ function usually used for pretraining image/text models via cross-entropy. Image/text pretraining allows the model to learn an aligned representation space for images and text; thus, adding more semantic information about the objects in an image to the model representations.
We use the SigLIP-So400m variant of SigLIP as our teacher model. The So400m variant uses an optimized ViT backbone whose performance is similar to one of the largest ViTs, ViT-g/14 [1] while having fewer parameters and thus being smaller. The number of parameters of SoViT-400m/14 lies somewhere between ViT-L/16 and ViT-g/14. The output dimensionality of the image and text encoder representations of SoViT-400m/14 is pp=$1152$ each. We align the image encoder representations with human odd-one-out choices using the UD optimization outlined in Eq. 6. This allows us to increase the triplet odd-one-out accuracy of SigLIP-So400m from $44.24\%$ to $61.7\%$ (see rightmost column in Tab. 1) which is close to the human-noise ceiling of $66.67\%$ for THINGS [cf. 28] and thus among the best human-aligned models without AligNet fine-tuning [cf. 55]. Note that this is a relative increase in performance of $39.47\%$. Throughout this manuscript we use the human-aligned version of SigLIP-So400m as the surrogate teacher model for generating human-like similarity judgments and distilling human-like similarity structure into student VFMs. We select a diverse and representative set of student VFMs.
@@ A.1.5 Student Models
Since previous research has demonstrated that a model’s architecture has no significant impact on the degree of alignment with human similarity judgments [55, 56], we use the same architecture for all student models that we fine-tune on AligNet. Specifically, we use the Vision Transformer [ViT; 17] for the backbone of each student model. We use the ViT rather than CNN-based model because ViTs have recently emerged as the dominant neural network architecture for computer vision application and VFMs. Every large VFM that is used in practice is based on the ViT [80, 13, 86, 6]. Unless otherwise mentioned, we use the base model size, i.e., ViT-B. ViT-B has $12$ attention layers and an internal (hidden) representation size of $p=768$. It has been shown that both the training data and the objective function have a substantial impact on the degree of alignment with human behavior. Thus, we use student models that were trained on different pretraining task with different training data and objective functions.
Supervised pre-training is still the prevailing mode of training computer vision models. Therefore, we trained ViT-B on the popular ImageNet dataset consisting of 1.4M natural images [16]. To examine how model performance changes as a function of the size of a model, we train ViT instances of three different sizes on ImageNet: ViT-S/16, ViT-B/16, ViT-L/16. The image patch size is the same for each of those models. To evaluate the effect of AligNet on self-supervised pretraining, we use pretrained DINOv1 [12] and DINOv2 [59] models of which DINOv1 was pretrained on ImageNet and DINOv2 was pretrained on a different, larger image dataset as denoted below. In addition, we investigate multimodal training of vision models that add textual information both in the form of image captioning via the CapPa model [80], as well as contrastive language-image pretraining (CLIP) via SigLIP [86]. The latter model is considered state-of-the-art on many downstream computer vision applications and is used as the image embedding model in modern large visual-language models [VLMs; 13, 6]. The full list of student models that we consider in our analyses is the following:
@@
    • Training data: ImageNet [16]
@@
  • CLIP (ViT-B) (see Sec.˜B.2.4)
    • Training data: WebImageText [66]
    • Objective: CLIP [66]
  • SigLIP (ViT-{B,SO400M})
    • Training data: WebLI [86]
    • Objective: CLIP [66]
  • SigLIP2 (ViT-B) (see Sec.˜B.2.4)
    • Training data: WebLI [86]
    • Objective: Combination of various objectives (see Tschannen et al. [81] for details)
@@
    • Training data: ImageNet [16]
    • Objective: Self-supervised image pretraining [12]
@@
    • Objective: Self-supervised teacher-student distillation [59]
@@
    • Objective: Multi-modal image captioning [80]
  • Randomly initialized ViT-B (see Sec.˜B.2.5)
    • Training data: AligNet
    • Objective: AligNet objective (see Eq.˜5 in Sec.˜A.1.3)
@@ A.2.1 Multi-arrangement task
Human similarity judgments for King et al. [37] and [14] were obtained by using a multi-arrangement task. In a multi-arrangement task, participants are presented with a computer screen showing images of several different objects. The participants are asked to arrange the images into semantically meaningful clusters, given the instruction that images of objects that lie close together are considered more similar. From this arrangement, one can infer pairwise (dis-)similarities of the objects and average those across all participants to obtain a representative (dis-)similarity matrix.
@@ A.2.3 Neural network representations
Representation Similarity Matrices (RSMs) for neural network representations are obtained by first embedding the same set of images that were presented to the human participants in the $p$-dimensional latent space of a model. The latent space could be any layer of a neural network. For the base models, we use the representations of the image encoder for SigLIP and the CLS token of the penultimate layer for CapPa, DINOv2, and ViT-B. We do this because previous work has shown that the penultimate layer space and the image encoder space of image/text models respectively yield the highest similarity to human behavior [63, 64, 55].
@@
[…]
@@ A.3 Levels data
Stimuli. The experimental stimuli were images taken from the ImageNet dataset [16]. Another 9 images were used for instructions only and depicted natural objects selected from the Bank of Standardized Stimuli [BOSS; 10], available at BOSS. We grouped the visual stimuli presented in the triplets according to different levels of abstraction: coarse-grained semantic, which comprised three images from three different categories; fine-grained semantic, showing three images from the same category; and class-boundary, where two images were from the same and one from a different category.
Instead of randomly sampling triplets—which would reproduce dataset biases—we stratified sampling by superclasses. ImageNet classes follow the WordNet hierarchy [16, 73], which includes higher-level classes. For instance, all dog breeds can be summarized as a single dog superclass. To avoid presenting dogs, birds, and other fine-grained classes that are overrepresented in ImageNet more frequently to the participants than other categories, we grouped the ImageNet classes into $717$ coarse-grained WordNet superclasses. We uniformly at random sampled images from those $717$ superclasses to construct the different kinds of triplets. Note that for all superclasses with more than one class, we uniformly at random chose one subclass and either uniformly at random sampled one image, two images (without replacement), or three images (without replacement) from that subclass, depending on the triplet type. For most superclasses that were comprised by a single subclass only, i.e., a one-to-one-mapping, we could skip the subclass sampling part. Triplet sampling resulted in ($N=450$) predefined experiment trial sets, of which ($N=448$) were used for testing. Across these, each triplet was presented within ($N=5$) different experiment files. This sampling process ensured a balanced distribution of triplets across the sample, and the repetition of each triplet in five different participants allowed for the calculation of an uncertainty distribution for each triplet.
@@
Procedure. The experiment was run online using jsPsych v7.3.3 and custom plugins. Participants were asked to provide demographic information, including their age and gender. Thereafter, they viewed written instructions about the task and performed six practice trials (2 trials per triplet level of abstraction). Participants were free to repeat the instructions until they felt confident to perform the experiment. The experiment proper comprised $N=330$ experiment trials. Each trial started with a fixation cross (1s), followed by the presentation of a triplet (max. 15s). Participants were asked to select the “odd-one-out” using the right, left, or downward facing arrow keys on their keyboard. Responses could be entered between 1-15s after triplet onset, after which the next trial started. Trials in which participants failed to submit a response were rare (M = 0.27% of trials; min = 0.00%, max = 6.06%). The serial order of triplet types (e.g., fine-grained or coarse-grained semantic) and ImageNet classes (e.g., dogs or birds) was counterbalanced across the experiment. We additionally counterbalanced the serial position of trial types across participants using a Latin-Square approach [LSD; 25]. Participants could take short breaks (self-paced for up to 2 minutes) after $N=50,150,$ and, $200$ experiment trials; Experimental trials were interleaved with $N=16$ catch trials (class border triplets), which were predefined based on low model uncertainty and 100% agreement among participants on these specific triplets during piloting. Catch trial performance was used as an indicator of adequate task engagement (see participant inclusion criteria above).
@@
Appendix B Additional/Detailed Results
In addition to the results that we presented in the main paper, here, we report the same set of results in more detail and add further analyses that we touched upon in the main part but did not go into much detail. We start by presenting findings from a comprehensive set of evaluations that we performed to test the alignment of model representations with human similarity judgments both for similarity judgments collected in previous efforts [cf. 63, 14, 37, 28] and for our own similarity judgment dataset Levels (see Sec. A.3). Subsequently, we demonstrate benefits of our method for machine learning downstream tasks that test generalization, such as few-shot learning and out-of-distribution detection. Finally, we perform a rigorous qualitative analysis of the changes in the model representations after applying our alignment framework to the various student models.
@@ B.1 Human Alignment
  • Distillation without alignment: We perform AligNet fine-tuning without the first step (see Fig.˜1) of aligning the teacher model with the human similarity judgments in THINGS. That is, we distill the teacher similarity structure of the pretrained, non-aligned SigLIP-So400m model (see Methods) into a student VFM.
@@ B.1.1 Detailed results for RSA
In Tab.˜1 we present the average alignment scores for all student models that we considered in our analyses (see Methods) and each of the four representation settings defined above for the datasets from Cichy et al. [14], King et al. [37], and Hebart et al. [28]. For the former two datasets, alignment is measured via RSA using Spearman rank correlation between the upper triangulars of the pairwise similarity matrices. For those datasets, human similarity judgments were collected using a multi-arrangment task where participants were asked to arrange (natural) objects on a computer screen (see Sec. A.2.1 for details). For the THINGS dataset, alignment is measured via triplet odd-one-out accuracy, i.e., the fraction of triplets for which models selected the same odd-one-out object as the human participants.
@@
ViT-L $0.041$ $\mathbf{0.540}$ $0.227$ $\mathbf{0.751}^{\dagger}$ $0.095$ $0.463$ $0.201$ $0.664$ $34.84\%$ $9.30\% 44$$4.40\%$ $60.75\%$
@@
DINOv2 ViT-B $\mathbf{0.492}$ $0.522$ $0.710$ $0.743$ $\mathbf{0.464}$ $0.445$ _\_ $0.661$ $\mathbf{51.24}\%$ $49.49\%$ 60.5160.51% $62.13\%$
@@
SigLIP So400m (Teacher) $0.226$ _ $0.723$ _ $0.267$ _ $\mathbf{0.638}$ _ $44.24\%$ _ $\mathbf{61.70}\%$ _
@@
We find that base models—that is the original representation setting (see above)—performed poorly for almost every human similarity judgment dataset (see Tab. 1 and Fig.˜6). The differences among the models in this setting is significant. While DINOv2 ViT-B achieves a Spearman rank correlation coefficient of $\rho=0.492,p<0.001$ for the dataset in Cichy et al. [14] and an odd-one-out accuracy of $51.21\%$ for the THINGS dataset, ViT-L is not correlated with the human similarity judgments from Cichy et al. [14] and King et al. [37] ($\rho=0.041$ and $\rho=0.095$ respectively) and shows a close to chance-level odd-one-out accuracy of $34.84\%$ for the THINGS dataset (see Tab. 1 and $x$-axis in Fig.˜7). AligNet fine-tuning significantly improved the alignment scores of all models for all datasets. In addition, it minimized the differences between the models to an extent that their differences in the degree of alignment are not statistically significant. Thus, AligNet fine-tuning models are equally well aligned with human similarity judgments irrespective of their architecture, pretraining data, and objective function.
In the top row of Fig.˜6 we show the performance of the transformed representations as a function of the base model performance for those datasets. We can see that AligNet fine-tuning significantly improves upon the base model representations ($x$-axis). It yields the most human-aligned models across all three datasets (depicted by the circles) and often substantially improves upon the linearly aligned representations (depicted by the squares).
@@
In the bottom two rows of Fig.˜6 and Fig.˜7 respectively we show model performances individually for the single category pairwise similarity rating datasets. We observe that AligNet fine-tuning significantly improved the degree of alignment for all models for the various, animals, and automobiles categories but there does not appear to be a difference in performance between the different model transformation settings for the categories fruits, furniture, and vegetables. We hypothesize that this difference in alignment benefits stems from the fact that animal- and automobile-related concepts are more frequently represented in the THINGS dataset than fruits-, furniture-, and vegetable-related concepts [cf. 27, 28]. Moreover, THINGS is a human similarity judgment dataset that reflects global coarse-grained semantic, similar to the coarse-grained setting in Levels. Thus, it is rather surprising to find benefits for human local fine-grained semantic by either UD or fine-tuning on AligNet.